### v1

# IndoCulture: Exploring Geographically-Influenced Cultural Commonsense Reasoning Across Eleven Indonesian Provinces

Abstract
Although commonsense reasoning is greatly shaped by cultural and geographical factors, previous studies on language models have predominantly centered on English cultures, potentially resulting in an Anglocentric bias. In this paper, we introduce IndoCulture, aimed at understanding the influence of geographical factors on language model reasoning ability, with a specific emphasis on the diverse cultures found within eleven Indonesian provinces. In contrast to prior works that relied on template Yin et al. (2022) and online scrapping Fung et al. (2024), we create IndoCulture by asking local people to manually develop the context and plausible options based on predefined topics. Evaluations of 23 language models reveal several insights: (1) even the best open-source model struggles with an accuracy of 53.2%, (2) models often provide more accurate predictions for specific provinces, such as Bali and West Java, and (3) the inclusion of location contexts enhances performance, especially in larger models like GPT-4, emphasizing the significance of geographical context in commonsense reasoning. [^1] IndoCulture can be accessed at https://huggingface.co/datasets/indolem/IndoCulture
1 Introduction
Figure 1: IndoCulture covers eleven provinces spanning from eastern to western Indonesia. The highlighted regions in the map represent the provinces examined in IndoCulture. We offer examples from Aceh, South Borneo, and Papua, with correct answers indicated in bold. English translations are for illustrative purposes.
Reasoning abilities of multilingual language models are frequently evaluated using English texts, potentially amplifying an Anglocentric bias toward English culture and leading to less inclusive models Thomas (1983); Ponti et al. (2020). Cultures, however, vary significantly from one location to another and profoundly shape the way a speaker of a language reasons Hershcovich et al. (2022). Recent evaluations of models reasoning ability OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Sengupta et al. (2023); Liu et al. (2023b) conducted on the English datasets, such as PIQA Bisk et al. (2020a), HellaSwag Zellers et al. (2019), Social IQA Sap et al. (2019), often overlook geographical aspects, thereby risking cultural bias.
Culture is a multifaceted concept encompassing the way of life Giddens and Sutton (2021), including our thoughts and actions Macionis John . It includes tangible elements like food art, and clothing, as well as intangible aspects such as ideas, values, attitudes, and norms. Culture is shaped by geographical location and ethnicity, influencing the commonsense reasoning of people within a region. For example, in Indonesia, it is culturally acceptable to eat rice with your hands but it is considered unusual to use a fork. Similarly, at traditional Indonesian weddings, it is common to sit on the floor while eating, whereas this practice is less common in Australia.
This work focuses on understanding the influence of geographical contexts in cultural commonsense reasoning, with the main focus on Indonesian culture. Indonesia is a highly multicultural country Putra et al. (2019), home to over 1,300 recognized ethnic groups and more than 700 languages Zarbaliyev (2017); Aji et al. (2022). As the the largest archipelagic country in the world, Indonesia has a population exceeding 270 million spread across 38 provinces, stretching from Aceh province in the west to Papua province in the east. Few prior studies on commonsense reasoning in Indonesian contexts Mahendra et al. (2021); Wibowo et al. (2023); Putri et al. (2024) overlooked the geographical nuance and the rich diversity of Indonesian cultures.
This paper introduces IndoCulture, a novel dataset to evaluate cultural reasoning in eleven Indonesian provinces, manually developed by local people in each province based on predefined topics. In prior works, cultural reasoning has primarily relied on datasets constructed through templates Yin et al. (2022), and online scraping Nguyen et al. (2023a); Fung et al. (2024). While these studies offer valuable insights, they may be susceptible to training data contamination when utilized to assess large language models (LLMs). For instance, Fung et al. (2024) reported a zero-shot accuracy of 92% when using ChatGPT to evaluate low-resource data.
Figure 1 illustrates the eleven provinces examined in IndoCulture including Aceh, North Sumatra, West Sumatra, West Java, Central Java, East Java, Bali, South Borneo, East Nusa Tenggara (NTT), South Sulawesi, and Papua. [^2] Although Papua consists of six provinces, for the purpose of this study, we treat it as a single entity (referred to as Papua) due to the relatively recent establishment of most of these provinces. These provinces span from the eastern to the western parts of Indonesia, each representing a major island in the country, with the addition of Bali and NTT. Figure 1 also shows three examples in IndoCulture for three provinces: Aceh, South Borneo, and Papua. The first two examples focus on cultural artifacts: the traditional wedding dress from Aceh and the Bagandut dance from South Borneo. The second example also examines cultural symbols. The third example focuses on cultural beliefs and norms regarding pregnancy in Papua.
Can large language models effectively reason based on the diverse cultures of Indonesia? To capture the rich diversity of Indonesian cultures, we meticulously predefined 12 fine-grained topics as guidelines for data construction. Figure 2 displays the topic distribution in IndoCulture, with the majority focusing on food, weddings, art, pregnancy and children, and family relationships. Additionally, we also pose the question: Is there an influence of geographical location on commonsense reasoning by language models? We address these questions through comprehensive experiments across different language models, incorporating several levels of location granularity as additional context in the prompt.
Our contributions can be summarized as follows:
  • We present IndoCulture, a high-quality cultural reasoning dataset in the Indonesian language, covering eleven provinces of Indonesia and twelve fine-grained cultural topics. Our dataset has 2,429 instances, and was meticulously developed by local people with rigorous quality controls in place.
  • We assess 15 open-source multilingual models, 6 open-source Indonesian-centric models, and 2 closed-source models. Although local individuals can answer all questions correctly (i.e., 100% accuracy), we observed that all open-source models struggle to comprehend Indonesian cultures, with the highest performance reaching only 53.2%.
  • We conduct a thorough analysis over various dimensions: (1) models performance for each province and topics, (2) the influence of different granularities of location contexts (i.e., none, province, country), (3) model performance over English translation, (4) analysis over model explanation behind its answer.
2 Related Work
Commonsense Reasoning in English
Many studies have focused on commonsense reasoning in English, often overlooking considerations of cultures and geographical locations. Early works include the Winograd Schema Challenge Levesque et al. (2012) and WinoGrande Sakaguchi et al. (2021) for pronoun coreference resolution tasks. Other research areas include reasoning based on cause-effect relationships Roemmele et al. (2011), physical activities Bisk et al. (2020b), social interaction Sap et al. (2019), cloze story completion Mostafazadeh et al. (2016), sentence completion Zellers et al. (2019), numerical Lin et al. (2020), and temporal Qin et al. (2021). Additionally, pretrained language models are employed in other works to extract structured commonsense knowledge by providing seed words Davison et al. (2019), and using code language model Madaan et al. (2022).
Cultural Commonsense Reasoning with Geographical Contexts
Previous studies have explored commonsense reasoning with geographical contexts. Shwartz (2022) investigated time perception (e.g., morning and night hours) across different locations, while Yin et al. (2022) delved into cultural knowledge of language models across five countries by masking and template strategies. Other works have focused on automatically extracting cultural knowledge from various sources, including Wikipedia Fung et al. (2024), conversations Fung et al. (2023), and Common Crawl Nguyen et al. (2023a), incorporating location contexts with the assistance of large language models. Related, Ziems et al. (2023) created a knowledge bank for situational norms, utilizing English-speaking Mechanical Turk annotators and incorporating country taxonomy. Unlike these works, IndoCulture specifically concentrates on cultural reasoning within Indonesian provinces, developed and validated manually by local people (experts). Compared to the automatic method and English-speaking crowdsource workers for data construction, IndoCulture arguably contains less noise, and is free from training data contamination of large language models (LLMs).
Commonsense Reasoning with Indonesian contexts
Commonsense reasoning in Indonesian language models has been studied using translated English–Indonesian datasets, such as XCOPA Ponti et al. (2020) and XStoryCloze Lin et al. (2021). However, these datasets potentially introduce a cultural bias toward English culture. IndoCloze Koto et al. (2022) is the first commonsense reasoning dataset in Indonesian, developed by native Indonesian workers following cloze story completion framework Mostafazadeh et al. (2016). IndoMMLU Koto et al. (2023), a benchmark for assessing knowledge in the Indonesian context, includes questions pertaining to Indonesian culture. However, its primary focus lies on testing proficiency in the local language rather than reasoning abilities. In a contemporary study, Wibowo et al. (2023) followed the COPA framework Roemmele et al. (2011) to build a dataset with contexts limited to Jakarta. In other work, Putri et al. (2024) studied the capability of LLMs in generating questions with cultural norms, for both general Indonesian and specific Sundanese contexts. In comparison, IndoCulture offers broader coverage across eleven provinces, thereby fostering greater inclusivity for local communities in Indonesia.
3 IndoCulture
As illustrated in Figure 1, IndoCulture is a sentence completion task in Indonesian language featuring a one-sentence premise, three plausible options, and one correct option to evaluate reasoning ability and cultural knowledge across eleven Indonesian provinces. In total, there are 2,429 instances, with the distribution across provinces detailed in Table 1.
3.1 Data Construction
IndoCulture is constructed manually by human and verified through a two-step process.
Worker Recruitment
Culture often arises from the shared experiences, traditions, and beliefs of a specific group over time, often closely intertwined with native populations. Keeping this in mind, we engaged individuals from various provinces across Indonesia to assist in preparing data for the IndoCulture benchmark. During the recruitment, we presented a few examples of the intended IndoCulture data and requested each candidate to generate similar instances tailored to the context of their respective provinces. From a pool of 58 applicants, we carefully selected 22 expert workers representing 11 provinces (with 2 workers selected per province). These recruited expert workers are local residents and have resided in their respective provinces for a minimum of 10 years, thereby possessing a profound understanding of local customs and culture. The age range of our workforce spans from 21 to 35 years old, with educational backgrounds distributed as follows: 3 high school graduates, 14 bachelor’s degree holders, 4 master’s degree holders, and 1 PhD holder. During data construction, each expert worker will fulfill the dual roles of instance writer and quality controller. [^3] Each worker is compensated above the monthly minimum wage in Indonesia.
Province Selection
The provinces covered in this study represent diversity of Indonesian cultures. Those 11 provinces (in Figure 1) spread across 6 islands and large archipelagos in Indonesia which are inhabited by various tribes who speak different regional languages and adhere to different religions.
Topic Taxonomy
IndoCulture consists of 12 topics and 66 fine-grained subtopics. Below is a list of topics along with their detailed subtopics. The numbers following the topic indicate the total instances required to be written by one worker (with a total of 150 per worker).
  1. 1.
    Food (22): breakfast (2); lunch (3); dinner (2); snacks (2); food souvenir (3); traditional foods and beverages (5); eating habit (1); cutlery (1); cooking ware (1), fruits (2).
  2. 2.
    Wedding (20): traditions before marriage (3); traditions when getting married (3); traditions after marriage (3); men’s wedding clothes (2); women’s wedding clothes (2); invited guests (2); wedding location (1); foods at a wedding (2); gifts brought to weddings (2).
  3. 3.
    Family relationship (13): relationships within the main family (3); relationships in the extended family (3); relations with society/neighbors (5); clan/descendant system (2).
  4. 4.
    Pregnancy and kids (16): traditions during pregnancy (4); traditions after birth (2); how to care for a newborn baby (2); how to care for toddlers (2); how to care for children (2); how to care for teenagers (2); parents and children interactions as adults (2).
  5. 5.
    Death (10): when death occurs (2); the process of caring for a corpse (2); traditions after the body is buried (2); the clothes of the mourners (2); inheritance matters (2).
  6. 6.
    Religious holiday (12): traditions before religious holidays (2); traditions leading up to religious holidays (4); traditions during religious holidays (5); traditions after religious holidays (1).
  7. 7.
    Agriculture (6): what to plant (2); traditions when planting (2); harvest (2).
  8. 8.
    Fisheries and trade (7): traditions of taking care of livestock/fish (5); buying and selling traditions (2)
  9. 9.
    Art (16): musical instruments (3); folk songs (3); traditional dances (3); use of art at certain events (5); poetry or similar literature (2)
  10. 10.
    Traditional games (5): game types; (3), location played (2).
  11. 11.
    Daily activities (10): morning activities (1); afternoon activities (1); evening activities (1); leisure activities (3); house, household and transportation (4).
  12. 12.
    Socio-religious aspects of life (13): regular religious activities (2); mystical things (2); traditional ceremonies (1); lifestyle (3); self care (1); traditional medicine (3); traditional saying (1).
Instance Writing
For each instance, workers craft two culturally relevant sentences that align with the predefined subtopic. The first sentence serves as the premise context, while the last sentence acts as the correct answer. Subsequently, the annotator generates two additional plausible sentences as distractors by modifying cultural objects or activities from the correct sentence. Workers are given a period of two months to complete the task.
Two Stages of Quality Control
In stage 1, we implement quality control by pairing two annotators from the same province. Each annotator is tasked with answering a set of questions prepared by the other annotator, and vice versa. During this phase, the annotator is presented with a premise sentence and three shuffled options. They are allowed to search for the answer from any source if they are unsure. Instances that are incorrectly answered by the second annotator are discarded, as we hypothesize that these instances may contain incorrect answers or possess a level of ambiguity. Additionally, annotators are required to identify whether the instance is province-specific (binary annotation: True/False), meaning it is uniquely relevant in their province and not in others.
In stage 2 of quality control, the first two authors of this paper perform post-editing on the data that has passed the first stage of quality control. We first focus on correcting the linguistic aspects of the text, which includes checking for spelling typos. Although the data is written in Indonesian, some annotators may use dialects or be influenced by the structure or style of regional languages. In these cases, we correct the text to adhere to correct Indonesian grammar.
To maintain the quality of IndoCulture, we rigorously filter instances that display (1) poor writing, as enhancing their quality substantially is difficult; (2) obvious answer options, which allow for easy guessing of the correct choice without understanding the cultural contexts; and (3) ambiguous contexts, where all options are equally valid as the correct answer. For example, in a topic about breakfast, the three options might include one traditional food alongside two other very commonly consumed foods in Indonesia.
Furthermore, we manually verify the province-specific annotation for each instance using the Google search engine. We annotate whether the instance pertains to national-level culture. If the example is specific to a province, it will be annotated as uncommon in national culture, and vice versa.
3.2 Data Statistics
After the annotation process, we initially collected 3,162 instances. [^4] Although we requested each annotator to produce 150 instances, not all were able to complete their allotted tasks within the given timeframe. Unfortunately, we were unable to find additional candidates from the same local province to address the data deficiencies Winata et al. (2023). In stage 1 of quality control, the initial pool of 3,162 instances was reduced to 2,801 instances. Our thorough stage 2 of quality control further reduced the sample to 2,429 final high-quality samples. The data distribution of IndoCulture per province is presented in Table 1. Approximately three-quarters of IndoCulture instances contain province-specific content, with an average length of around 35 words. IndoCulture covers multiple topics, as illustrated in Figure 2.
Figure 2: Topic distribution in IndoCulture.
Province # province $\mu$(word) $\mu$(char)
specific (%)
Aceh 246 70.7 28.0 175.9
North Sumatra 234 83.8 36.8 246.0
West Sumatra 299 74.6 39.6 261.4
West Java 231 58.0 37.5 244.8
Central Java 171 66.7 39.3 260.5
East Java 233 69.5 46.0 310.4
Bali 241 76.3 33.3 216.1
NTT 103 72.8 31.8 203.6
South Borneo 233 83.7 33.3 226.0
South Sulawesi 185 90.3 33.6 227.8
Papua 253 88.1 37.3 245.0
All 2429 76.0 NA NA
Table 1: Overall statistics of IndoCulture by province.
4 Experiments
4.1 Set-Up
We focus on evaluating 23 language models in zero-shot settings: (1) 15 open-source multilingual language models including BLOOMZ Muennighoff et al. (2022), mT0 Muennighoff et al. (2022), LlaMA2 Touvron et al. (2023), Bactrian-X Li et al. (2023a) in different sizes, (2) 2 South East Asian language models including SeaLLM Nguyen et al. (2023b), and SeaLion Singapore (2023), (3) 4 Indonesian-centric language models, including IndoBART Cahyawijaya et al. (2021), IndoGPT Cahyawijaya et al. (2021), Merak Ichsan (2023), Komodo Owen et al. (2024), (3) 2 closed-source models, including ChatGPT: gpt-3.5-turbo Ouyang et al. (2022) and GPT–4: gpt-4-0613 OpenAI (2023). Please refer to Appendix A for further details.
Initially, we evaluate the effectiveness of sentence completion and multiple-choice question strategies in predicting the correct options. In both scenarios, we conduct benchmarks across three distinct location contexts. Formally, given a premise $s$, three candidate options $c_{1}$, $c_{2}$, $c_{3}$, and location $l\in\{\texttt{none},\texttt{Indonesia},p_{j}\}$, for sentence completion, we select the correct option based on:
[…]
Here, $\text{concat}(s,c)$ denotes the concatenation of premise $s$ and candidate option $c$, separated by a space. In the case of multiple-choice questions, we devise a template for the prompt question and determine the answer by selecting the option with the highest probability among letters A, B, and C. Figure 3 displays the prompt templates in Indonesian and English for both sentence completion and multiple-choice question.
Figure 3: Templates for sentence completion and multiple-choice questions prompts.
4.2 Results
Model (#parameter) Completion MCQ
$l=\text{None}$ $l=\text{Ind}$ $l=\text{Prov}$ $l=\text{None}$ $l=\text{Ind}$ $l=\text{Prov}$
Human 100.0 100.0
Random 33.3 33.3 33.3 33.3 33.3 33.3
BLOOMZ (560M) 37.5 35.3 35.3 32.1 32.0 32.1
BLOOMZ (1.1B) 36.2 36.9 37.2 32.0 32.0 32.0
BLOOMZ (3B) 38.5 40.5 41.2 45.9 47.6 48.6
BLOOMZ (7B) 40.9 43.8 44.4 48.8 49.8 49.7
[…]
LLaMA2 (7B) 36.7 37.0 37.2 40.0 39.0 37.7
LLaMA2-chat (7B) 36.9 36.9 37.4 40.5 41.1 40.3
LLaMA2 (13B) 39.4 39.7 39.9 46.9 46.9 46.6
LLaMA2-chat (13B) 38.2 38.4 38.9 47.2 49.2 49.2
[…]
IndoBART (132M) 41.9 41.1 41.7 32.1 32.0 32.2
IndoGPT (117M) 42.6 42.1 42.7 33.7 34.1 35.0
Merak (7B) 40.2 40.6 42.6 51.5 52.5 52.5
SeaLLM (7B) 39.1 39.7 41.1 51.7 52.4 52.5
SEA-LION (7B) 38.8 38.7 39.6 33.4 32.8 33.0
Komodo (7B) 45.1 45.4 46.1 37.6 35.1 36.1
GPT–3.5 (NA) 58.4 60.2 61.7
GPT–4 (NA) 68.8 71.5 75.8
Table 2: Zero-shot accuracy across various models and settings. “MCQ” refers to the multiple-choice question method, and $l$ denotes the location as additional context. The bold numbers highlight the highest score within each model group.
Overall observation
[…]
Multiple-choice question method is generally better.
Our findings suggest that the multiple-choice question method tends to outperform the sentence completion method, with exceptions noted for BLOOMZ (560M, 1.1B), IndoBART, IndoGPT, and Komodo. Interestingly, within the sentence completion method, the small Indonesian-centric language model, IndoGPT, surpasses the performance of almost all large multilingual models and remains competitive with BLOOMZ (7B). However, IndoGPT’s performance experiences a significant decline in the multiple-choice question method, with a notable margin of nearly 8 to 10 points. While Komodo (7B) achieves the highest accuracy with the sentence completion method, its performance also decreases by 10-12 points in the multiple-choice method. This discrepancy could potentially be attributed to differences in the nature of language model training and instruction-tuning.
Impact of location context on model performance
Our investigation reveals that incorporating various levels of location granularity has a noticeable effect on zero-shot performances, especially models with larger parameter sizes. Detailed location contexts notably enhance the accuracy of BLOOMZ (7B), LlaMA2 (13B), Merak (7B), SeaLLM (7B), Komodo (7B), GPT–3.5, and GPT–4. For instance, in GPT–4, the accuracy gap between $l=\texttt{none}$ and $l=\texttt{Indonesia}$ is 2.7, and this gap further increases by 7 points when $l=\texttt{Prov}$ is assigned.
4.3 Analysis
Given the superior performance of models with large parameter sizes using the multiple-choice question method and location $l=\texttt{Prov}$, we employ these settings for analysis.
Province Merak GPT–3.5 GPT–4
$\neg$¬PS PS $\neg$¬PS PS $\neg$¬PS PS
Aceh 59.7 53.4 79.2 65.5 93.1 73.6
North Sumatra 58.3 46.9 63.9 57.2 75.0 73.2
West Sumatra 50.0 41.3 63.2 48.4 85.5 63.7
West Java 65.3 58.8 84.2 71.8 92.6 81.7
Cental Java 58.8 47.1 67.6 52.9 82.4 72.9
East Java 57.4 37.6 72.2 46.5 87.0 63.4
Bali 78.9 65.6 87.7 81.7 93.0 89.4
NTT 64.3 52.0 57.1 49.3 85.7 68.0
South Borneo 65.7 46.4 68.6 53.6 77.1 67.4
South Sulawesi 53.3 47.9 66.7 55.8 80.0 70.6
Papua 76.7 55.2 76.7 62.8 90.0 71.3
Table 3: Top-3 model accuracy by province. "PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
Results by province
Table 3 highlights that the top 3 performing LLMs exhibit a nuanced understanding of cultures within Indonesian provinces, particularly excelling in the cultures of West Java and Bali compared to other provinces. GPT–4, for instance, achieves the best accuracy of 93%, while Merak and GPT–3.5 also demonstrate strong performance in Bali and West Java, with accuracies ranging between 55% and 88%. In other provinces like West Sumatra and South Borneo, the models typically exhibit poorer performance. Specifically, for GPT-3.5, the performance gap compared to Bali ranges from 19 to 33%. This highlights the presence of cultural biases and a lack of inclusivity in model reasoning abilities, likely stemming from the distribution of training data. The proximity of West Java to Jakarta (Indonesia’s capital) and Bali’s global status as a tourism destination may contribute to the abundance of textual data on these two cultures.
We also note a consistent disparity between non-province and province-specific contexts across all models, with models generally finding non-province contexts easier to comprehend. On average, this gap ranges from 12 to 13 points for the three models, highlighting the challenge posed by province-specific content and emphasizing the significant influence of location context on the reasoning ability of LLMs.
Results by topic
Table 4 shows the accuracy of the top 3 performing models across topics. Similar to Table 3, the models perform better in province-specific context for all topics, with death customs being an exception. For province-specific contexts, GPT–4 excels in themes of wedding, death, and arts while for non-specific contexts, GPT–4 achieves accuracy more than 90.0 for family relationship, religious holidays, and farm and fishery.
Province Merak GPT–3.5 GPT–4
$\neg$¬PS PS $\neg$¬PS PS $\neg$¬PS PS
Food 56.9 51.0 65.9 62.1 79.3 63.5
Wedding 66.7 64.3 77.6 56.0 92.8 78.6
Family relationship 65.5 54.5 72.5 53.9 93.1 72.7
Pregnancy and kids 67.9 42.9 69.2 50.6 82.1 72.7
Death 38.5 55.2 60.9 62.5 76.9 82.5
Religious holidays 68.5 48.8 79.3 68.8 94.4 64.6
Agriculture 70.5 52.3 71.4 62.1 88.5 76.1
Farm and fishery 77.8 51.0 77.8 51.9 94.4 75.8
Art 58.7 50.2 85.2 61.1 78.3 77.3
Traditional game 60.0 50.2 80.3 59.8 88.0 66.9
Daily activities 55.1 39.6 66.0 64.8 87.8 63.0
Socio-religious 57.1 44.3 79.6 61.1 85.7 64.6
Table 4: Top-3 model accuracy by topic. "PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
Does language affect model performance?
We automatically translated IndoCulture to English using the Google Translate API [^5] Accessed in March 2024. and used the English prompt in Figure 3 to evaluate the models. All model performances in English text dropped except for LLaMA-2 and Merak. This could be attributed to two reasons. First, LLaMA-2 is an English-centric model, and Merak is pretrained from LLaMA-2. Second, the performance drop in other models could be caused by translation errors. We further investigated this with 100 random samples and found that 81 samples had acceptable translations. We observed translation errors such as pronoun mismatches, inaccurate proverb translations, and inaccurate translations of local terms, such as “pupuik” translated as “fertilizer”.666pupuik is a traditional musical instrument in West Sumatra. It is worth noting that the word pupuik closely resembles the word pupuk in Indonesian, which means fertilizer. To better understand the cultural gap in language models, we followed the approach of Liu et al. (2023a) to manually correct the translations and reevaluate the models. We found that GPT-4’s performance over the 100 random samples was 77.0 for the original Indonesian text, 68.0 for the English machine-translated text, and 72.0 for the English translation fixed by humans.
[…]
Results by fine-grained culture elements
We tasked two expert workers with annotating 200 random samples from IndoCulture based on six cultural elements, as derived from Axtell and Fornwald (1998); Williams (2014). [^7] https://www.languageeducatorsassemble.com/elements-of-culture [^8] https://pressbooks.howardcc.edu/soci101/chapter/3-2-the-elements-of-culture/ While these elements may not encompass every cultural aspect, we contend that they cover the most prominent or pivotal elements, including: (1) symbols (material or non-material objects representing meaning), (2) artifacts (material or non-material objects produced by society), (3) values and beliefs (principles, ideas, and concepts assumed to be ideal and correct in society), (4) norms (rules guiding values and beliefs), (5) language, (6) rituals (established procedures and ceremonies), and (7) other, for examples that do not fit into any of the defined elements. This annotation is a multi-label task, and the average Kappa score across the cultural elements is 0.56, with each ranging from 0.4 to 0.75. These scores indicate moderate to substantial agreement.
Table 5 displays the distribution of each cultural element in our dataset, along with the performance breakdown across Merak, GPT-3.5, and GPT-4. Among the 200 random samples, we observe that 42.5% of our data contains artifacts, 37.5% norms, and 30% rituals. Only 4% of the data pertains to symbols, while 7.5% belongs to the other category. Merak exhibits lower performance in norms, with a 24% decrease compared to values and beliefs. Conversely, GPT-3.5 and GPT-4 perform better in norms, with the performance gap in values and beliefs being only $-4$ points. Additionally, language poses a challenge in Merak, with only 38% accuracy, while GPT-3.5 and GPT-4 achieve 60% and 72% accuracy, respectively.
Cultural element (%) Merak GPT–3.5 GPT-4
Symbols (4) 50.0 58.9 70.8
Artefacts (42.5) 55.3 59.1 67.8
Values and Beliefs (10.5) 61.9 57.5 69.8
Norms (37.5) 38.7 63.2 73.6
Language (19.5) 38.5 60.2 72.0
Ritual (30) 53.3 60.0 70.7
Other (7.5) 66.7 58.9 69.2
Table 5: Accuracy comparison of Merak, GPT–3.5, and GPT–4 across 200 random samples, categorized by cultural elements. The numerical value following each cultural element indicates its proportion within the samples.
Can the model provide a reasonable explanation to support the answer?
Figure 5: Performance comparison between Merak (7B) and GPT–4 based on text generation output. “Answer (T)” indicates that the generated answer is true, while ”Exp(F)” denotes that the answer explanation is false.
We conduct a manual investigation of the text generation output for Merak (7B) and GPT–4 across 200 random samples. This involves manually examining the generated answer along with its explanation. To obtain explanation, we modify the Indonesian prompt in Figure 3 by adding the string Jelaskan jawabanmu! ("Explain your answer!"). Our annotation process is binary, categorizing explanations as either True or False. We label an explanation as False if it is absent, contains hallucinations, or provides inaccurate information. [^9] We use Google search engine to verify the correctness of the explanation.
As anticipated, there is a significant drop in accuracy for Merak (7B) from 52.5% (as shown in Table 2) to 29.5%. This discrepancy underscores the limitations of relying solely on token probability to assess the true capability of a language model. Interestingly, only 4.5% of the samples are answered correctly with the appropriate explanation by Merak, despite it being the top performer among the open-source models. GPT–4 achieves an accuracy of 69.5%, which is 5 points lower than indicated in Table 2, possibly due to a smaller sample size and biases in regular expressions. However, GPT–4 provides better explanations, with 51.5% of the samples correctly answered along with the appropriate explanation.
5 Discussion
The recent study Wang et al. (2024) has demonstrated that evaluating language models using token probability in multiple-choice question types does not align well with the generated text. This discrepancy is reported to be more pronounced in models fine-tuned on conversational or safety data. In response to this issue, we have conducted a manual evaluation on 200 random samples (in Section 4.3) and found that the performance in generated texts, especially for Merak, the best open-source language model, is deteriorating. However, this issue is less apparent in GPT-4, as the initial evaluation is based on regular expressions. Conducting manual evaluation on all data and models is expensive, and we plan to address this issue in future works. This work primarily focuses on introducing a novel dataset meticulously constructed for evaluating cultural commonsense reasoning within the Indonesian context, including preliminary evaluation results based on the standard methods used in previous studies OpenAI (2023); Touvron et al. (2023); Koto et al. (2023); Li et al. (2023b); Koto et al. (2024).
6 Conclusion
IndoCulture is a cultural commonsense reasoning dataset encompassing the diversity of Indonesian cultures, spanning from Aceh province in the west to Papua province in the east. Through collaboration with local individuals across eleven provinces and rigorous quality control measures, we introduce IndoCulture for the purpose of evaluating language models. Our findings reveal that large language models, whether Indonesian-centric or multilingual, demonstrate a limited understanding of Indonesian culture. Notably, incorporating location as additional context significantly enhances model performance, particularly in GPT–3.5 and GPT–4.
Limitations
IndoCulture is specifically designed to explore the influence of geographical location on cultural commonsense reasoning, with a focus on the present time. It does not consider temporal aspects. Our dataset was created in the year 2023, recognizing that cultures may evolve over time, as discussed by Mesoudi (2016).
Furthermore, as demonstrated in Section 4.3, a significant portion of IndoCulture comprises cultural elements such as artifacts, norms, and rituals. Symbols, values and beliefs, and language represent smaller proportions, ranging from 4% to 20%. We encourage future research to further explore these cultural elements and to expand the geographical coverage beyond the eleven provinces studied in this paper.
Ethical Considerations
IndoCulture is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. [^10] https://creativecommons.org/licenses/by-nc-sa/4.0/ Our data is intended for academic research and non-commercial purposes. Workers are compensated above the minimum monthly salary in Indonesia and are fully aware that the data will be released to the public. It is important to note that no private or sensitive information of the workers is included in IndoCulture.
Appendix A Model Artifacts
Models (#parameters) Source
BLOOMZ (560M) bigscience/bloomz-560m
BLOOMZ (1.1B) bigscience/bloomz-1b1
BLOOMZ (1.7B) bigscience/bloomz-1b7
BLOOMZ (3B) bigscience/bloomz-3b
BLOOMZ (7.1B) bigscience/bloomz-7b1
[…]
LLamA2-chat (7B) meta-llama/Llama-2-7b-chat
LLamA2 (13B) meta-llama/Llama-2-13b
LLamA2-chat (13B) meta-llama/Llama-2-13b-chat
[…]
IndoGPT (117M) indobenchmark/indogpt
Merak (7B) Ichsan2895/Merak-7B-v5-PROTOTYPE1
SeaLLM (7B) SeaLLMs/SeaLLM-7B-v2
SEA-LION (7B) aisingapore/sea-lion-7b
Komodo (7B) Yellow-AI-NLP/komodo-7b-base
Table 6: With the exception of GPT-3.5 and GPT-4, all the models used in this study were sourced from Huggingface Wolf et al. (2020).